General Meeting Summary 3/22/17

From cua_phy
Jump to navigation Jump to search

JLAB DETECTORS

  • Discussion about SHMS Aerogel Cherenkov detector (Arthur, Hamlet, Vardan)
  • Position dependence data from KPP runs have been analyzed. Results (horizontal and vertical in the spectrometer coordinate system, where X is pointing down and Y beam right) show negligible dependence on vertical position (note and no dependence on horizontal position
  • Discussion about getting SP-11 tray ready
  • Walter's group is constructing the cover plate
  • Next steps: 1) install last layer of SP-11 aerogel tiles and support, 2) install wires, 3) install tray in SHMS and perform cosmics tests
  • PMT cabling optimization to be done before next hall closing in Fall
  • for possible options keep in mind that cannot have negative impact on tray exchanges/sliding out detector from detector stack and work on other detectors near aerogel
  • NPS prototype (Arthur, Hamlet)
  • Linearity tests of HV dividers have been completed with single crystal prototype - results show that stable up to 1-2 MHz consistent with design specs
  • Discussion about tests of NPS cooling systems with 3x3 prototype - idea is to use a closed-circuit design, need to get a suitable chiller
  • Gluex DIRC (Greg)
  • Discussion about transportation of DIRC bars from Stanford Linear Accelerator to JLab
  • bars will be in a crate and transported by truck, monitoring cameras will be installed for all twelve bars, reference spectra will be monitored during transportation, lead vehicle to warn about uneven roads
  • discussion about adding a trailing vehicle to address rear-ending


PIXEL DETECTORS (Rachel)

  • Discussion about shipment to CUA
  • on schedule to arrive on 3/24
  • discussion about procurement of an optical table for gantry
  • Discussion about detectors at CERN
  • good progress - sealed, dry, cooled - calibration data have been taken in control room


PBWO4 STUDIES AT CUA (Richard, Andres)

  • working on assembly of the prototype, e.g. cutting aluminum to size, drilling holes etc (Richard)
  • working on attentuator circuit to reduce reflections in APD readout (Andres)


TIMELIKE COMPTON SCATTERING (Vardan, Arthur, Pawel)

  • Work on preliminary projections and comparison to already approved experiments ongoing
  • Proposal deadline is 22 May, but also need to circulate to NPS collaboration
  • Decision on whether or not to move forward with PAC proposal this year will depend on outcome of preliminary projections


NEXT MEETING: WED. 29 March at 10:00 AM (EDT)